CURRENT RESEARCH ON HYDROGEN-BONDED COMPLEXES AT THE UNIVERSITY OF VAL-LADOLID

\underline{A. LESARRI}, M. E. SANZ, S. ANTOLÍNEZ, S. BLANCO, J. C. LÓPEZ and J. L. ALONSO, Departamento de Química Física, Facultad de Ciencias, Universidad de Valladolid, 47005 Valladolid, Spain.

Investigation of hydrogen-bonded heterodimers formed by the six-membered ring heterocycles tetrahydropyran or thiane with HCl has recently evidenced the formation of both axial and equatorial hydrogen bonds^{a} ^{b}.

We have extended these studies to other complexes in which the proton acceptor molecule of the complex has nonequivalent binding sites. In this communication we present an analysis of the conformational preferences and structures of the hydrogen-bonded complexes thietane-HCl, methylenecyclobutane-HCl and cyclopentene oxide-HCl, investigated by molecular-beam Fourier transform microwave spectroscopy (MB-FTMW).

aS. Antolínez, J. C. López and J. L. Alonso, Angew. Chem., in press (1999)

bM. E. Sanz, J. C. López and J. L. Alonso, submitted (1999)